The following sets have been defined using the | notation. Re-write them by listing some of the elements.


i. {p | p is a capital city, p is in Europe}


ii. {z | 3z = n2, z and n are natural numbers}



1
Expert's answer
2022-06-13T17:45:47-0400

i. {p | p is a capital city, p is in Europe}={Athens, Berlin, Copenhagen, Dublin, London, Madrid, Paris, Oslo, Riga, Rome, Vienna, ...}


ii. {"z | 3z = n^2," z and n are natural numbers}

We have: "z=\frac{n^2}{3}."

As z and n are natural numbers, n is divizible by 3.

At n=3: "z=\frac{3^2}{3}=3,"

at n=6: "z=\frac{6^2}{3}=12,"

at n=9: "z=\frac{9^2}{3}=27,"

at n=12: "z=\frac{12^2}{3}=48, ..."

So

{"z | 3z = n^2," z and n are natural numbers} = {3, 12, 27, 48, ...}